OpenStudy (anonymous):

What is 1,000,000 written as a power of ten?

OpenStudy (anonymous):

10^9

OpenStudy (anonymous):

or is it 10^6? because in scientific notation, you'll find that 1,000,000 is 1 x 10^6. Count the zeros

OpenStudy (anonymous):

Sorry, you are right. I miscounted on accident. It is 10^6.

OpenStudy (whpalmer4):

\[10^2 = 10*10 = 100\]\[10^3 = 10*10*10 = 1,000\]\[10^4=10*10*10*10 = 10,000\]\[10^5= 10*10*10*10*10=100,000\]\[10^6=10*10*10*10*10*10=1,000,000\]